A hollow rectangular waveguide has dimensions a = 2.28 cm., b = 1.01 cm. What TE modes will propagate in this waveguide if the frequency of the electromagnetic waves is 1.7 x 1010 Hz? What range of frequencies can be used to excite only one TE mode?


1
Expert's answer
2020-08-04T16:01:13-0400

As per the given question,

Dimensions of the wave guide "a = 2.28 cm, b = 1.01 cm"

Frequency of the electromagnetic wave "(f)=1.7\times 10^{10}Hz"

Range of frequencies =?

Cut off wave guide for the dominant "TE_{10}" mode "=\frac{2}{\sqrt{(\frac{m}{a})^2+(\frac{n}{b})^2}}"

"=\frac{2}{\sqrt{(\frac{1}{2.28})^2+(\frac{0}{1.01})^2}}"

"=0.046 cm"

For the higher wavelength "=\frac{2}{\sqrt{(\frac{m}{a})^2+(\frac{n}{b})^2}}"

"=\frac{2}{\sqrt{(\frac{1}{2.28})^2+(\frac{1}{1.01})^2}}"

"=4.02 cm"

Hence the required frequencies

"f_c =\frac{c}{2a} =\frac{3\times 10^8}{2\times 2.28 \times 10^{-2}}Hz"

"=0.657 \times 10^{10}Hz"

and

"f_c=\frac{c}{2b}"


"=\frac{3\times 10^8}{2\times 1.10 \times 10^{-2}}Hz"


"=1.36\times 10^{10}Hz"